OpenStudy (anonymous):

Find the sum of the infinite geometric sequence. 1 -1/5 + 1/25 - 1/125 . . .

OpenStudy (anonymous):



OpenStudy (anonymous):

here first term = 1 common ratio = -1/5 no. of terms is infinite

OpenStudy (anonymous):

the general formula for sum of Geometric Series till infinity is a/(1-r) plug in the values

OpenStudy (anonymous):

wouldn't it be 1/1-1/5 = 1/ 2/5

OpenStudy (anonymous):

no